# Teaching Transformers Causal Reasoning through Axiomatic Training

### v2

@@ D.1 Definitions
Appendix E Positional Encodings and their Role in Generalization
Positional Encoding (PE) play a crucial role of providing information about the absolute and relative position of tokens in a sequence (Vaswani et al., 2023). (Vaswani et al., 2023) propose an absolute positional encoding strategy using periodic functions (e.g., sinusoidal or cosine) to initialize these encodings. Absolute positional encoding provides definite values for all positions across any sequence length. However, studies (Ontañón et al., 2022; Csordás et al., 2021) show absolute positional encoding fails in length generalization tasks for transformers. In the learnable APE variant (Radford et al., 2018), each positional embedding is randomly initialized and trained with the model. This approach falters with sequences longer than those seen in training, as the new positional embeddings remain untrained and randomized. Interestingly, recent findings (Kazemnejad et al., 2023; Haviv et al., 2022) indicate that removal of PEs in auto-regressive models can improve model’s length generalization capabilities, wherein the attention mechanism during auto-regressive decoding is sufficient to encode positional information. We also experiment with Rotary Position Encodings (Su et al., 2023), which have shown superior length generalization. We use $\theta=10000.0$ for the base period of RoPE embeddings.
Appendix F Formalising training and Evaluation Setup
Let $f_{dim}$ represent the maximum value for a given perturbation dimension $dim$, along which we construct train and evaluation sets for our axiomatic framework. For each dimension, we choose a threshold $\tau_{dim}\in L$, such that $f_{dim}<\tau_{dim}$ forms our training set and $f_{dim}\geq\tau_{dim}$ forms the evaluation set. So, $f_{dim}\in\{f_{len},\ f_{branch},\ f_{nodelen},\ f_{revfactor},\ f_{shuffle}\}$ where:
@@ Appendix F Formalising training and Evaluation Setup
[…]
@@
Appendix G Results of Dsep on CLEAR dataset
@@
{mdframed}
[linecolor=black,linewidth=0.5pt]       Premise: Given a DAG (directed acyclic graph) with nodes C, Z, P, V and directed edges C->V, P->V, C->Z, Z->P, Z->V.
Hypothesis: "Which of the following nodesets can d-separate node C and node P?
@@
Figure A1: Example instance of Multiple Choice (MC) question type from (Chen et al., 2024b) dataset describing d-separation rule problem defined with a different hypothesis type and semantic structure then the one our models are finetuned on.
Appendix H Results on Corr2Cause dataset
Figure A2: Model Comparison: F1 Score across Templates, finetuning on dsep and transitivity based axiomatic instances lead to performance improvement. Model finetuned on transitivity sees the highest jump for templates like Parent, where identifying direct-indirect relations is important. Finetuning on D-separation instances see a jump in performances for templates like has_collider, where identifying a collider is important. Since child template has 0 ’No’ labels, the F1 score reported is 0.
Figure A3: Comparison of Model Performance across Different Relationship Templates. Accuracy plots show consistent trends, with model finetuned on transitivity consistently outperforming base models, on templates where direct-indirect relationship identification is required. Finetuning on D-separation instances see a performance jump over the base instruct model, for templates identifying colliders and confounders.
Appendix I Implementation Details
We used a learning rate of 1e-4 with linear scheduling and 3% warmup ratio, training for 4102 max steps on axiomatic instance samples with sequences of maximum length 4096 tokens. We employed mixed precision (bfloat16) training with flash attention for efficiency. After training, the LoRA weights were merged with the base model for inference. We used Huggingface (wol, 2020) for implementation. The fine-tuning used LoRA with rank 64, alpha 16, and dropout 0.1. Training was performed on 3 GPUs using DeepSpeed Stage 3 with a total batch size of 128 (16 samples per GPU with gradient accumulation). 1 A40 and 1 A100 GPUs were used for training the transformer model from scratch for all Positional encodings based experiments.

### v3

@@ E.1 Definitions
Appendix F Positional Encodings and their Role in Generalization
Positional Encoding (PE) play a crucial role of providing information about the absolute and relative position of tokens in a sequence (Vaswani et al., 2017). (Vaswani et al., 2017) propose an absolute positional encoding strategy using periodic functions (e.g., sinusoidal or cosine) to initialize these encodings. Absolute positional encoding provides definite values for all positions across any sequence length. However, studies (Ontañón et al., 2022; Csordás et al., 2021) show absolute positional encoding fails in length generalization tasks for transformers. In the learnable APE variant (Radford et al., 2018), each positional embedding is randomly initialized and trained with the model. This approach falters with sequences longer than those seen in training, as the new positional embeddings remain untrained and randomized. Interestingly, recent findings (Kazemnejad et al., 2023; Haviv et al., 2022) indicate that removal of PEs in auto-regressive models can improve model’s length generalization capabilities, wherein the attention mechanism during auto-regressive decoding is sufficient to encode positional information. We also experiment with Rotary Position Encodings (Su et al., 2024), which have shown superior length generalization. We use $\theta=10000.0$ for the base period of RoPE embeddings.
Appendix G Formalising training and Evaluation Setup
Let $f_{dim}$ represent the maximum value for a given perturbation dimension dimdim, along which we construct train and evaluation sets for our axiomatic framework. For each dimension, we choose a threshold $\tau_{dim}\in L$, such that $f_{dim}<\tau_{dim}$ forms our training set and $f_{dim}\geq\tau_{dim}$ forms the evaluation set. So, $f_{dim}\in\{f_{len},\ f_{branch},\ f_{nodelen},\ f_{revfactor},\ f_{shuffle}\}$ where:
@@ Appendix G Formalising training and Evaluation Setup
  • Let $l_{i,j}$ be the length of the name of the node $X_{i,j}$, then $l_{i,j}=(len(X_{i,j}))$.Therefore, the maximum length of node names across all nodes in all causal sequences can be represented as: $f_{nodenamelen}=\max_{1\leq i\leq n,\,1\leq j\leq m}l_{i,j}$. We set $\tau_{nodelen}$ for train set as 3, with $f_{nodelen}\in[1,3]$.
@@
Appendix H Results of Dsep on Longer Causal Chains With Randomly Flipped Edges
@@
Appendix I Examples of instances from each benchmark: CLEAR and Corr2Cause
Premise: Given a DAG (directed acyclic graph) with nodes C, Z, P, V and directed edges C->V, P->V, C->Z, Z->P, Z->V. Hypothesis: "Which of the following nodesets can d-separate node C and node P?
@@ Appendix I Examples of instances from each benchmark: CLEAR and Corr2Cause
Figure A1: Example instance of Multiple Choice (MC) question type from (Chen et al., 2024) dataset describing d-separation rule problem defined with a different hypothesis type and semantic structure than the one our models are fine-tuned on.
Premise: Suppose there is a closed system of 4 variables, A, B, C and D. All the statistical relations among these 4 variables are as follows: A correlates with B. A correlates with C. A correlates with D. B correlates with C. B correlates with D. C correlates with D. However, B and D are independent given A. B and D are independent given A and C. C and D are independent given A. C and D are independent given A and B. Hypothesis: There exists at least one collider (i.e., common effect) of A and B. Label: No.
Figure A2: Example instance from the Corr2Cause dataset, where the model must infer the presence of a collider between variables given only correlational and conditional independence statements.
Appendix J Implementation Details
We used a learning rate of 1e-4 with linear scheduling and 3% warmup ratio, training for 4102 max steps on axiomatic instance samples with sequences of maximum length 4096 tokens. We employed mixed precision (bfloat16) training with flash attention for efficiency. After training, the LoRA weights were merged with the base model for inference. We used Huggingface (Wolf et al., 2020) for implementation. The fine-tuning used LoRA with rank 64, alpha 16, and dropout 0.1. Training was performed on 3 GPUs using DeepSpeed Stage 3 with a total batch size of 128 (16 samples per GPU with gradient accumulation). One A100 GPU and one A40 GPU were used for training the transformer model from scratch for all Positional Encodings based experiments, and 4 A6000 GPUs were used for finetuning experiments.